# Boundary crisis and long transients of the Atlantic overturning circulation mediated by an edge state

### v1

@@ Abstract
The Atlantic Meridional Overturning Circulation (AMOC), a large-scale ocean current and vital climate component, could transition to a substantially weakened flow state. Despite severe associated impacts, it remains difficult to reliably estimate the proximity to potential critical thresholds and predict the AMOC’s fate under global warming. To advance our understanding of Earth system stability, a global view on the dynamics is needed beyond the local stability analysis underlying, e.g., standard early warning signs. Here we explore the phase space of an intermediate-complexity climate model featuring a metastable AMOC. For two atmospheric carbon dioxide (CO2) levels, we explicitly compute the edge state (also called Melancholia state), a chaotic saddle embedded in the basin boundary that separates the strong and weak AMOC attractors found in the model. While being unstable, the edge state can govern the transient climate for centuries. Its dynamics exhibits strong centennial AMOC oscillations driven by atmosphere-ice-ocean interactions in the North Atlantic. We demonstrate that at CO2 levels projected to be reached in the next decades, the simulated AMOC undergoes a boundary crisis where the current AMOC attractor collides with the edge state. Near but beyond the crisis, long chaotic transients occur due to a resulting ghost state, causing a splitting of ensemble trajectories under time-dependent forcing. Rooted in dynamical systems theory, our results offer an interpretation of large ensemble variances and a so-called stochastic bifurcation observed in earth system models under intermediate forcing scenarios.
@@ I Introduction
Earth’s climate is a metastable complex system [1]: on various scales, the variability of paleoclimate records is characterised by relatively abrupt transitions between distinct long-lived climatic regimes [2, 3, 4]. From a dynamical systems perspective, we may interpret the observed metastability by regarding the Earth system as a forced multistable system featuring a hierarchy of competing attracting states [5, 6]. The stability landscape of the underlying autonomous system is thereby described by a global quasipotential based on Graham’s field theory [7, 8], with local minima of the landscape corresponding to attractors.
@@
The quasipotential landscape of the Earth system has been explored in the context of the Cenozoic Era [9] and for our planet’s bistability in the extent of glaciation (“Snowball Earth”) [10, 11, 6]. Here, we close in on a subscale feature of the present-day climate thought to be multistable: the Atlantic Meridional Overturning Circulation (AMOC), a widely studied system of large-scale ocean currents [12, 13, 14]. The AMOC plays a vital role in climate by transporting heat northwards, supporting northern Europe’s relatively mild climate [15]. A suspected driver of past climate metastability [12], the AMOC is one of the proposed climate tipping elements – Earth system components susceptible to abrupt and possibly irreversible large-scale change [16, 17]. Given the ongoing anthropogenic climate change [18], there is growing concern that one or more tipping elements could cross a tipping point and transition to a qualitatively different state, with severe consequences for humanity and nature [19, 20]. The possibility of tipping events complicates climate prediction, contributing large uncertainty to risk management and adaptation strategies.
@@
How the AMOC will respond to global warming is an urgent open question. Climate models forced with emissions scenarios until the year 2100 consistently project a decline of the AMOC, though its magnitude is model-dependent [21, 18]. While some studies infer a recent weakening from reconstructions [22], direct measurements are short and noisy [23]. According to the latest assessment of the Intergovernmental Panel on Climate Change (IPCC AR6), a collapse before 2100 is “unlikely (medium confidence)” but cannot be ruled out [18]. An AMOC shutdown would have severe global impacts, including a relative cooling of the northern hemisphere, reduction in precipitation and increased storminess in Europe, shifts of rainfall patterns globally, and regional accelerations in sea level rise [24, 25]. Even without a full shutdown, a partial collapse of the circulation could result from the shutdown of deep convection zones in the North Atlantic Subpolar Gyre (SPG) [26, 27]. Such a transition could cause qualitatively similar impacts within decades [28].
@@
A hierarchy of climate models – from box models [29] to intermediate-complexity models [30, 31] and comprehensive earth system models [32, 33] – indicates that the AMOC can be multistable [34, 35, 36]. In a certain regime of atmospheric heat and freshwater forcing, a vigorous flow state resembling today’s circulation (ON state) coexists with a much weaker or collapsed overturning state (OFF state). While there could be additional competing states [37], this bistability underlies the classical view of the AMOC as a tipping element. The bistability stems from the positive salt-advection feedback, describing the interdependence between the AMOC flow strength and northward salt transport [38, 15], which could be triggered by surface buoyancy changes in the North Atlantic.
@@
To address the risk of an AMOC transition, research has aimed at detecting early warning signs (EWS), determining critical forcing levels, and estimating tipping times. A series of recent studies [39, 40, 41] has applied statistical EWS to time series of observed AMOC reconstructions, suggesting that the AMOC is approaching a tipping point that could be reached this century [41]. These methods rely on the concept of critical slowing-down as the system approaches a bifurcation. An indicator based on the salt import into the Atlantic supported these findings when applied to reanalysis data [42]. However, these methods have serious limitations that make the robust prediction of a transition difficult in practice, if not prohibitive [43]. A crucial underlying assumption is that the system remains close to an equilibrium state, which may not hold given the current rate of anthropogenic forcing. Instead, the scenario of non-autonomous or rate-dependent tipping [17, 44] is more appropriate, for which an EWS theory is missing and indicators based on critical slowing-down fail [45].
@@
The trajectory of the Earth system constitutes one realization of a chaotic complex system [46]. This implies limits to predictability intrinsically linked with chaos [47, 48]. Near critical thresholds of metastable systems, the sensitive dependence on the initial condition (predictability of the first kind [47]) can strongly inhibit the predictability of the asymptotic state (second kind) [49, 50]. Particularly, an ensemble of trajectories may partially tip under identical time-dependent forcing, simply due to internal variability: some ensemble members transition, while others do not. Ensemble splitting has been found in climate models of intermediate and high complexity [51, 52, 50, 53]. In the NASA GISS-E2-1-G Earth system model (hereafter GISS model) of the Coupled Model Intercomparison Project Phase 6 (CMIP6), an ensemble sampled from internal variability showed divergent AMOC behaviour described as a “stochastic bifurcation” [52]. By construction, critical slowing-down indicators cannot discern between such trajectories that tip and those that do not [45]. To understand this behaviour and predict tipping in chaotic non-autonomous systems, a global stability view beyond stable equilibria is needed.
@@
While studies often emphasise the binary question of tipping or not tipping, the transient behaviour can be equally important [54, 55]. Long transients and metastable dynamics are often governed by unstable states (non-attracting invariant sets), which are underexplored in climate models [56]. A particular class of unstable states are edge states, also called Melancholia states [10, 57]: saddles embedded in the basin boundaries that partition the phase space between the competing stable states. Edge states may be defined as attractors of the dynamics restricted to the basin boundary – pictorially, “mountain passes” between valleys of the global quasipotential landscape [58]. Thus, edge states commonly act as gateways of critical transitions (with caveats [59]). Numerically, edge states can be found using an edge tracking algorithm originally proposed in the context of turbulent flow [60, 61, 62] and recently applied to climate models [10, 48, 57, 58].
@@
In complex systems, the basin boundaries may exhibit a highly complicated geometry [48, 63, 10]. Rather than marking a sharp threshold, the boundary can be a gray zone in phase space of fractal dimension where the system is virtually unpredictable. In parameter space, this translates to a fuzzy tipping region, as opposed to a clear tipping point, when an attracting state collides with the boundary in a so-called boundary crisis [55, 64]. In the tipping region, the system may undergo long transients, such that a transition might occur thousands of years after a loss of stability. These aspects highlight that determining a precise tipping threshold or timing may not be possible, requiring a more holistic assessment of stability.
@@
In this work, we take a global view on the phase space of the coupled climate system. We explore the stability landscape of PlaSim-LSG, an Earth system model of intermediate complexity featuring a bistable AMOC under present-day conditions. Instead of focusing on the climates of the stable AMOC states, we investigate the edge state that lies in between and its role for transient dynamics. Building on recent studies using a conceptual climate model [48] and a global ocean circulation model [57], the key novelty here is that we consider a fully coupled modelling setup with a simplified yet Earth-like description of the ocean, atmosphere, cryosphere, hydrosphere and biosphere and their interactions.
@@
We perform edge tracking at two CO2 levels representing the recent and possible future climate, respectively. In the bistable regime (see Fig. 1), the edge state found in the model exhibits strong AMOC oscillations on centennial timescales, revealing a much richer dynamics compared with Ref. [57] driven by atmosphere-ice-ocean interactions. Combining simulations under autonomous and non-autonomous forcing, we demonstrate that in our model the AMOC undergoes a boundary crisis at CO2 levels exceeded under intermediate and high greenhouse gas emissions scenarios proposed by the IPCC. At the crisis, the ON state merges with the edge state giving rise to a so-called ghost state, a long-lived yet unstable chaotic set [65, 48, 66]. Near but beyond the crisis, we observe centennial to millennial transient behaviour which alternates between modes of variability reminiscent of the ON and edge states, before the circulation ultimately approaches the OFF state. Our findings help explaining the key aspects of the divergent AMOC behaviour as observed in the GISS model [52].
@@
Our paper is structured as follows. After introducing the model, we describe its AMOC bistability for the present-day climate and assess the AMOC response to future CO2 forcing scenarios until the year 3000 (section II). Then, we implement the edge tracking algorithm to construct an edge state of the AMOC (section III) and characterise its dynamical and physical properties (section IV). In section V, we explore the morphed stability landscape at a CO2 level beyond the boundary crisis, allowing to interpret the dynamics observed under time-dependent CO2 forcing (section VI).
@@ II.1 Model configuration
The atmosphere component of the Planet Simulator (PlaSim) [68] solves the moist primitive equations, describing the conservation of mass and momentum as well as basic thermodynamics, using simplified parameterisations of radiation, convection, precipitation and cloud processes. The prognostic equations are formulated in a spectral represenation truncated at T21 resolution horizontally (roughly giving a grid) with 10 vertical levels. The atmosphere is coupled to the Large Scale Geostrophic (LSG) ocean model [69] via a thick slab ocean serving as the surface layer of LSG that regulates air-sea fluxes. Assuming that the nonlinear terms of the Navier-Stokes equations can be neglected for large-scale ocean flows [80], the model solves the equations for momentum, temperature and salinity based on hydrostatic balance and the Boussinesq approximation. Convection is not explicitly resolved but accounted for via a convective adjustment scheme. At each time step, the scheme mixes vertically adjacent grid boxes whenever they are unstably stratified, starting from the top and iterating through the water column. Discretised on an E grid [81], LSG has an effective horizontal resolution of and 22 vertical layers on a stretched grid with thicknesses ranging from at the surface to in the deep ocean. The thermodynamic sea ice module is based on a zero-layer model [82] that computes the ice thickness from the thermodynamic balance at the ice-air and ice-ocean interface, accounting for snowfall. Sea ice transport is neglected.
We configure the model to roughly reflect present-day climatic conditions (with orbital parameters corresponding to around 2000 CE). Its climate sensitivity lies just above [70], consistent with the CMIP6 range [83]. At the baseline atmospheric CO2 concentration of (a level recorded in 1995), the default initialization of the model produces an AMOC with a volume transport of around at 26∘N (), close to today’s observed value of [84]. Here the AMOC strength is defined as the maximum of the Atlantic meridional streamfunction at a given latitude , taken over the depth (below sea level), with
@@ II.2 Transient CO2 forcing experiments
As a motivating experiment, we force the model with CO2 projections of Shared Socioeconomic Pathways (SSPs), standardised scenarios for greenhouse gas concentrations until 2500 [85] (Fig. 2a). For each SSP, we launch an ensemble of simulations (10 members) starting in the forcing year 1995, with initial conditions branched off from a 2000-year control run at . We compare low emissions (SSP1-2.6), intermediate emissions (SSP2-4.5), and high emissions (SSP4-6.0). Beyond 2500, we assume that the CO2 concentration decays exponentially to at the rate reached in the decade before 2500.
The AMOC shows qualitatively different behaviour under the three climate change scenarios (Fig 2b). For SSP1-2.6, the vigorous AMOC state is maintained over the 1000-year simulation, as exemplified for three ensemble members. In the SSP4-6.0 scenario, the AMOC collapses in the North Atlantic for all ensemble members. The abrupt decline starts after 2100 and happens within a century. Strikingly, in the intermediate SSP2-4.5 scenario, the ensemble splits, with the AMOC at 46-66∘N sometimes persisting and sometimes collapsing after strongly varying transients. Even though all ensemble members experience an identical time-dependent forcing, the internal variability leads to a qualitatively differing AMOC response. This difference imprints itself on the global climate, including the global mean surface temperature (Fig. 2c). Global warming under SSP2-4.5 can differ by up to depending on the state of the AMOC. Generally, AMOC weakening reduces the global mean surface temperature, in line with expectations [86].
@@ II.3 Bistability of the AMOC
At CO2, the model features (at least) two distinct stable AMOC states: a strong overturning cell with an average strength of (ON state) and a much weaker and shallower overturning circulation that shuts down to less than about north of 46∘N (OFF, Figs. 3a, b). Their stability has been verified via 4000-year long unforced simulations. Determining the precise CO2 range of the bistable regime is challenging due to the occurrence of long transients, as we discuss below.
@@
Still, the OFF state is characterised by the typical climate signal associated with an AMOC collapse, including a reduction of mean surface air temperature in the Northern Hemisphere (locally exceeding ), a drying of North Atlantic regions including northern Europe, and a southward shift of the tropical rain belt (Intertropical Convergence Zone) (Fig. 3) [24, 25]. We also find a strengthened polar jet stream in the northern hemisphere, combined with a large-scale reduction of zonal winds in other regions.
The time-averaged sea surface temperature (SST) is more than 2∘C (up to 9∘C) colder compared to the ON state in large parts of the North Atlantic, while the Southern Ocean is up to 3∘C warmer. The Atlantic subtropical gyre, southern Atlantic, Indian Ocean and Southern Ocean are saltier in the OFF state, whereas the North Atlantic and Arctic Ocean are substantially fresher (with the exception of the IrmS). This is a clear signature of the salt-advection feedback and meridional ocean heat transport: the weakened AMOC transports less salinity and heat from the tropics to the north.
The AMOC is closely connected with sites of deep oceanic convection in the North Atlantic, where dense water sinks. In models and observations, major deep convection sites are located in the Labrador Sea (LabS), Irminger Sea (IrmS), and Norwegian Sea (NorS; see Fig. S1 of the Supplemental Information 222The Supplemental Information is attached at the end of this preprint. for a map). In PlaSim-LSG, the transition from the ON to the OFF state is characterised by a shutdown of deep convection in the LabS and NorS (Fig. 3f), while the convection depth (as defined in the Supplemental Information) increases in several other locations.
In summary, the AMOC ON and OFF states have a qualitatively different climate on a global scale. In order to understand the transition behaviour between these states, we now investigate the global stability of the AMOC beyond the steady states.

### v2

# Global stability of the Atlantic overturning circulation: Edge states, long transients and a boundary crisis under CO2 forcing

@@ Abstract
The Atlantic Meridional Overturning Circulation (AMOC), a crucial ocean current system, could transition to a weak state. Despite severe associated climate impacts, assessing the AMOC’s response under global warming and its proximity to possible critical thresholds remains difficult. To understand future Earth system stability, a global dynamical view is needed beyond the local stability analysis underlying classical early-warning methods. Using an intermediate-complexity climate model, we explore the stability landscape of the AMOC for different atmospheric CO2 concentrations. We explicitly compute the edge state (or Melancholia state), a chaotic saddle on the basin boundary separating the strong and weak AMOC attractors found in the model. While being unstable, the edge state can govern the transient climate for centuries, supporting centennial AMOC oscillations driven by atmosphere-ice-ocean interactions in the North Atlantic. At increased CO2 levels projected for the near future, we reveal a boundary crisis where the current AMOC attractor disappears by colliding with the edge state. Under crisis overshoot, long chaotic transients due to “ghost states” lead to diverging ensemble trajectories under time-varying forcing. Rooted in dynamical systems theory, our results offer an explanation of large ensemble variance and apparent “stochastic bifurcations” observed in earth system models under intermediate forcing scenarios.
@@ I Introduction
Earth’s climate is a metastable complex system [1]: on various scales, the variability of paleoclimate records is characterised by relatively abrupt transitions between distinct long-lived climatic regimes [2, 3, 4]. From a dynamical systems perspective, we may interpret the observed metastability by regarding the Earth system as a forced multistable system featuring a hierarchy of competing attracting states [5, 6]. The stability landscape of the underlying time-frozen system is thereby described by a global quasipotential based on Graham’s field theory [7, 8], with local minima of the landscape corresponding to attractors.
@@
The quasipotential landscape of the Earth system has been explored in the context of the Cenozoic Era [9] and for our planet’s multistable extent of glaciation (“Snowball Earth”) [10, 11, 6]. Here, we close in on a subscale feature of the present-day climate thought to be multistable: the Atlantic Meridional Overturning Circulation (AMOC), a widely studied system of large-scale ocean currents [12, 13, 14]. The AMOC plays a vital role in climate by transporting heat northwards, supporting northern Europe’s relatively mild climate [15]. A suspected driver of past climate metastability [12], the AMOC is one of the proposed climate tipping elements [16, 17]. Given the ongoing anthropogenic climate change [18], there is growing concern that one or more tipping elements could cross a tipping point and transition to a qualitatively different state, with severe consequences for humanity and nature [19, 20]. The possibility of tipping events complicates climate prediction, contributing large uncertainty to risk management and adaptation strategies.
How the AMOC will respond to global warming is an urgent open question. Climate models forced with greehnouse gas emissions scenarios until the year 2100 consistently project a decline of the AMOC, though its magnitude is model-dependent [21, 18]. While some studies infer a recent weakening from reconstructions [22], direct measurements are short and noisy [23].
@@
An AMOC shutdown would have severe global impacts, including a relative cooling of the northern hemisphere, reduction in precipitation and increased storminess in Europe, shifts of rainfall patterns globally, and regional accelerations in sea level rise [24, 25]. Even without a full shutdown, a partial collapse of the circulation could result from the shutdown of deep convection zones in the North Atlantic Subpolar Gyre (SPG) [26, 27]. Such a transition could cause qualitatively similar impacts within decades [28].
@@
A hierarchy of climate models – from box models [29] to intermediate-complexity [30, 31] and comprehensive earth system models [32, 33] – indicates that the AMOC can be multistable [34, 35, 36]. In a certain regime of atmospheric heat and freshwater forcing, a vigorous flow state resembling today’s circulation (ON state) coexists with a much weaker or collapsed overturning state (OFF state). While there could be additional competing states [37], this bistability underlies the classical view of the AMOC as a tipping element. The bistability stems from the positive salt-advection feedback, describing the interdependence between the AMOC flow strength and northward salt transport [38, 15], which could be triggered by surface buoyancy changes in the North Atlantic.
@@
To address the risk of an AMOC transition, research has aimed at detecting early warning signs (EWS), determining critical forcing levels, and estimating tipping times. A series of recent studies [39, 40, 41] has applied statistical EWS to time series of observed AMOC reconstructions, suggesting that the AMOC is approaching a tipping point that could be reached this century [41]. These methods rely on the concept of critical slowing-down as the system approaches a bifurcation. An indicator based on the salt import into the Atlantic supported these findings when applied to reanalysis data [42]. However, these methods have serious limitations that make the robust prediction of a transition difficult in practice, if not prohibitive [43]. A crucial underlying assumption is that the system remains close to an equilibrium state, which may not hold given the current rate of anthropogenic forcing. Instead, the scenario of nonautonomous or rate-dependent tipping [17, 44] is more appropriate, for which an EWS theory is missing and indicators based on critical slowing-down fail [45].
@@
The trajectory of the Earth system constitutes one realization of a chaotic complex system [46]. This implies limits to predictability intrinsically linked with chaos [47, 48]. Near critical thresholds of metastable systems, the sensitive dependence on the initial condition (predictability of the first kind [47]) can strongly inhibit the predictability of the asymptotic state (second kind) [49, 50]. Particularly, an ensemble of trajectories may partially tip under identical time-dependent forcing, simply due to internal variability: some ensemble members transition, while others do not. Ensemble splitting has been found in climate models of intermediate and high complexity [51, 52, 50, 53]. In the NASA GISS-E2-1-G Earth system model (hereafter GISS model) of the Coupled Model Intercomparison Project Phase 6 (CMIP6), an ensemble sampled from internal variability showed divergent AMOC behaviour described as a “stochastic bifurcation” [52]. By construction, critical slowing-down indicators cannot discern between such trajectories that tip and those that do not [45]. To understand this behaviour and predict tipping in chaotic nonautonomous systems, a global stability view beyond stable equilibria is needed.
@@
While studies often emphasise the binary question of tipping or not tipping, the transient behaviour can be equally important [54, 55]. Long transients and metastable dynamics are often governed by unstable states (non-attracting invariant sets), which are underexplored in climate models [56]. A particular class of unstable states are edge states, also called Melancholia states [10, 57]: saddles embedded in the basin boundaries that partition the state space between the competing stable states. Edge states may be defined as attractors of the dynamics restricted to the basin boundary – pictorially, “mountain passes” between valleys of the global quasipotential landscape [58]. Thus, edge states often act as gateways of critical transitions (with caveats [59]). Numerically, edge states can be found using an edge tracking algorithm originally proposed in the context of turbulent flows [60, 61, 62] and recently applied to climate models [10, 48, 57, 58].
@@
In complex systems, the basin boundaries may exhibit a highly complicated geometry [48, 63, 10]. Rather than marking a sharp threshold, the boundary can be a gray zone in state space of fractal dimension where the system is virtually unpredictable. In parameter space, this leads to a tipping window, as opposed to a sharp tipping point, initiated when an attracting state collides with the boundary in a so-called boundary crisis [55, 64]. In the tipping window, the system may undergo long transients, such that a transition might occur thousands of years after a loss of stability [65]. These aspects highlight that determining a precise tipping threshold or timing may not be possible in finite time, requiring a more holistic assessment of stability.
@@
In this work, we take a global view on the state space of the climate system. We explore the stability landscape of PlaSim-LSG, an Earth system model of intermediate complexity featuring a bistable AMOC under present-day conditions. Instead of focusing on the climates of the stable AMOC states, we investigate the edge state that lies in between and its role for transient dynamics. Building on recent studies using a conceptual climate model [48] and a global ocean circulation model [57], the key novelty here is that we consider a fully coupled modelling setup with a simplified yet Earth-like description of the ocean, atmosphere, cryosphere, hydrosphere and their interactions.
@@
We perform edge tracking at three different CO2 concentrations representing the preindustrial, present-day and possible future climate. In the bistable regime (see Fig. 1), the edge state found in the model exhibits strong AMOC oscillations on centennial timescales, revealing a much richer dynamics than seen in Ref. [57], driven by atmosphere-ice-ocean interactions. Combining simulations under autonomous and nonautonomous forcing, we demonstrate that in our model the AMOC undergoes a boundary crisis at CO2 levels exceeded even under intermediate emission scenarios proposed by the IPCC. At the crisis, the ON state merges with the edge state giving rise to a so-called ghost state, a long-lived yet unstable chaotic set [66, 48, 67]. Near but beyond the crisis, we observe centennial to millennial transient behaviour which alternates between modes of variability reminiscent of the ON and edge states, before the circulation ultimately approaches the OFF state. Our findings help explaining the key aspects of the divergent AMOC behaviour observed in more comprehensive earth system models [52].
@@
Our paper is structured as follows. After introducing the model, we describe its AMOC bistability for the present-day climate and assess the AMOC response to transient CO2 forcing until the year 3000 CE (section II). In section IV, we implement the edge tracking algorithm to construct an edge state of the AMOC (section III) and characterise its dynamical and physical properties . In section V, we explore how the stability landscape changes as a function of CO2 level, revealing a boundary crisis. Relating this to transient simulations in a reduced state space allows to interpret the dynamics observed under time-dependent CO2 forcing (section VI, where we directly compare with the GISS simulations [52]).
@@ II.1 Model configuration
The atmosphere component of the Planet Simulator (PlaSim) [69] solves the moist primitive equations, describing the conservation of mass and momentum as well as basic thermodynamics, using simplified parameterisations of radiation, convection, precipitation and cloud processes. The prognostic equations are formulated in a spectral represenation truncated at T21 resolution horizontally (roughly giving a grid) with 10 vertical levels. The atmosphere is coupled to the Large Scale Geostrophic (LSG) ocean model [70], whereby the 50 m-thick uppermost vertical layer of LSG is used to compute air-sea fluxes. Assuming that the nonlinear terms of the Navier-Stokes equations can be neglected for large-scale ocean flows [81], the model solves the equations for momentum, temperature and salinity based on hydrostatic balance and the Boussinesq approximation. Convection is not explicitly resolved but accounted for via a convective adjustment scheme. At each time step, the scheme mixes vertically adjacent grid boxes whenever they are unstably stratified, starting from the top and iterating through the water column. Discretised on an E grid [82], LSG has an effective horizontal resolution of and 22 vertical layers on a stretched grid with thicknesses ranging from 50 m at the surface to 1000 m in the deep ocean. The thermodynamic sea ice module is based on a zero-layer model [83] that computes the ice thickness from the thermodynamic balance at the ice-air and ice-ocean interface, accounting for snowfall. Sea ice transport is neglected.
We configure the model to roughly reflect present-day climatic conditions (with orbital parameters corresponding to around 2000 CE). Its climate sensitivity lies just above C [71], consistent with the CMIP6 range [84]. At the baseline atmospheric CO2 concentration of 360 ppm (a level recorded in 1995), the default initialization of the model produces an AMOC with a volume transport of around 16 Sv at 26∘N ( m3 s-1), close to today’s observed value of Sv [85]. Here the AMOC strength is defined as the maximum of the Atlantic meridional streamfunction at a given latitude , taken over the depth (below sea level), with
@@ II.2 Transient CO2 forcing experiments
As a motivating experiment, we force the model with CO2 projections of Shared Socioeconomic Pathways (SSPs), standardised scenarios for greenhouse gas concentrations until 2500 [86] (Fig. 2a). For each SSP, we launch an ensemble of simulations (10 members) starting in the forcing year 1995, with initial conditions branched off from a 2000-year control run at 360 ppm. We compare low emissions (SSP1-2.6), intermediate emissions (SSP2-4.5), and high emissions (SSP4-6.0). Beyond 2500, we assume that the CO2 concentration decays exponentially to 330 ppm at the rate reached in the decade before 2500.
The AMOC shows qualitatively different behaviour under the three climate change scenarios (Fig 2b). For SSP1-2.6, the vigorous AMOC state is maintained over the 1000-year simulation, as exemplified for three ensemble members. In the SSP4-6.0 scenario, the AMOC collapses in the North Atlantic for all ensemble members. The abrupt decline starts after 2100 and happens within a century. Strikingly, in the intermediate SSP2-4.5 scenario, the ensemble splits, with the AMOC at 46-66∘N sometimes persisting and sometimes collapsing after strongly varying transients. Even though all ensemble members experience an identical time-dependent forcing, the internal variability leads to a qualitatively differing AMOC response. This difference imprints itself on the global climate, including the global mean surface temperature (Fig. 2c). Global warming under SSP2-4.5 can differ by up to 1∘C depending on the state of the AMOC. Generally, AMOC weakening reduces the global mean surface temperature, in line with expectations [87].
@@ II.3 Bistability of the AMOC
At 360 ppm CO2, the model features (at least) two distinct stable AMOC states: a strong overturning cell with an average strength of 16 Sv (ON state) and a much weaker and shallower overturning circulation that shuts down to less than about 2 Sv north of 46∘N (OFF, Figs. 3a, b). Their stability has been verified via 4000-year long unforced simulations. Determining the precise CO2 range of the bistable regime is challenging due to the occurrence of long transients, as we discuss below.
@@
Still, the OFF state is characterised by the typical climate signal associated with an AMOC collapse, including a reduction of mean surface air temperature in the Northern Hemisphere (locally exceeding 10∘C), a drying of North Atlantic regions including northern Europe, and a southward shift of the tropical rain belt (Intertropical Convergence Zone) (Fig. 3) [24, 25]. We also find a strengthened polar jet stream in the northern hemisphere, combined with a large-scale reduction of zonal winds in other regions.
The time-averaged sea surface temperature (SST) is more than 2∘C (up to 9∘C) colder compared to the ON state in large parts of the North Atlantic, while the Southern Ocean is up to 3∘C warmer. The Atlantic subtropical gyre, southern Atlantic, Indian Ocean and Southern Ocean are saltier in the OFF state, whereas the North Atlantic and Arctic Ocean are substantially fresher (except the Irminger Sea). This is a clear signature of the salt-advection feedback and meridional ocean heat transport: the weakened AMOC transports less salinity and heat from the tropics to the north.
The AMOC is closely connected with sites of deep oceanic convection in the North Atlantic, where dense water sinks. In models and observations, major deep convection sites are located in the Labrador Sea (LabS), Irminger Sea (IrmS), and Norwegian Sea (NorS; see Fig. S1 of the Supplemental Information for a map). In PlaSim-LSG, the transition from the ON to the OFF state is characterised by a shutdown of deep convection in the LabS and NorS (Fig. 3f), while the convection depth (as defined in the Supplemental Information) increases in several other locations.
In summary, the AMOC ON and OFF states have a qualitatively different climate on a global scale. To understand the transition behaviour between these states, we now investigate the global stability of the AMOC beyond the steady states.